# LAPS-Diff: A Diffusion-Based Framework for Singing Voice Synthesis With Language Aware Prosody-Style Guided Learning

### v1

@@ III-D Inference Stage
In the inference stage, the optimal auxiliary decoder $AD^{*}(\cdot)$ generates the mel-spectrogram $\hat{M}$, which contains nuanced representations of the lyrical content, vocal style, and prosody. The pre-trained HiFi-GAN vocoder [31] is then used to reconstruct the audio waveform from $\hat{M}$, from which the embeddings $\boldsymbol{e}_{{mert}}$ and $\boldsymbol{e}_{{vec}}$ are extracted. These embeddings along with $\boldsymbol{e}_{{c}}$, $\boldsymbol{e}_{{t}}$, and the intermediate boundary-predicted [1] representation $\hat{M}_{q}$ are used to condition the optimal denoiser $D^{*}(\cdot)$, as illustrated in Fig. 2 to generate the final output. Here, $\hat{M}_{q}$ is derived from $\hat{M}$ through the shallow diffusion mechanism of DiffSinger [1], where $q$ denotes the shallow diffusion step. Boundary prediction mechanism is applied only at the first iteration of the reverse diffusion, rest are same as Eq.(III-C).
@@
11:Compute reconstruction loss: $\mathcal{L}_{\text{L1}}=\text{MAE}(M,\hat{M})$
@@
29:   $\hat{M}_{q}=\text{ShallowDiffusion}(\hat{M},\epsilon),\quad\epsilon\sim\mathcal{N}(0,I)$
@@
40:        $M_{t-1}={}&\frac{1}{\sqrt{\alpha_{t}}}\Bigg{(}M_{t}-\frac{1-\alpha_{t}}{\sqrt{1-\bar{\alpha}_{t}}}\hat{\epsilon}_{\theta}(M_{t},\boldsymbol{e}_{c},\boldsymbol{e}_{t},\\ &\boldsymbol{e}_{\text{mert}},\boldsymbol{e}_{\text{vec}})\Bigg{)}+\sigma_{t}z\end{aligned}$
@@ IV-A Data Preprocessing
The overall duration of the singing voice segments of our Hindi Bollywood SVS dataset, introduced in Section II, is approximately 65 minutes. With 38 unique songs, segmented into 397 sung phrases each with duration ranging from 5 to 15 seconds. The dataset splits are based on distributing the songs across training, validation, and test sets with details shown in Table I. Each audio file is sampled at $16$ kHz with $16$-bit quantization. The Hindi text (Devanagari script) is converted into its corresponding phonemes using the IIT-M Hindi phoneset [32]. The size of the phoneme vocabulary considered is $43$. We extract mel-spectrograms considering $80$ mel-bins using frame size of $512$ and hop size of $128$. Similar to DiffSinger [1], the mel-spectrograms are normalized to the range $[-1,1]$.
@@ IV-A1 Implementation Details
In our work, we use the IndicBERT model [15], pre-trained on $12$ Indian languages, including Hindi. In contrast, the pre-trained XPhoneBERT model [16], which follows the BERT-base architecture [27], is trained on nearly $100$ languages, also including Hindi. The embedding dimension for both IndicBERT and XPhoneBERT is $768$. Linear projection is applied to reduce the dimensionality to 256, ensuring compatibility with the $256$-dimensional music score embedding $\boldsymbol{e}_{m}$. We retain the encoder architecture $E(.)$ as used in DiffSinger [30].
@@
The style encoder $S(.)$ consists of four 2D residual blocks [18], and the style vector dimensionality is set to 48. We use the checkpoint of the pre-trained JDCNet (a ConvBi-LSTM model consisting of three ResNet blocks) [19] from StyleTTS2 [17] to extract pitch values (the pitch sequence is of variable length depending on the segment duration). Unlike the loss functions described in the original DiffSinger paper [1], the official GitHub implementation [30] incorporates three additional duration losses—word, phone, and slur—duration along with a pitch and energy loss, employing duration, pitch and energy predictors similar to FastSpeech 2 [33]. These additional losses are also used to train the auxiliary decoder of both DiffSinger and our proposed model in all the experiments. The model architecture of $AD(.)$ is the same as FastSpeech2 [34, 1]). In LAPS-Diff, the denoiser architecture $D(.)$ follows a non-causal WaveNet framework [35], similar to that used in DiffSinger [1].
@@
The MERT model considered in our work is trained on diverse datasets comprising musical instruments and singing voices [20]. We use the pre-trained model to obtain the musical embeddings of size 1024. The IndicWav2Vec [21] is trained on nine Indian languages including Hindi, is capable of extracting linguistic content from speech audio. We employ the pre-trained IndicWav2Vec to obtain content embeddings of size 1024. We consider batch size as 48 and optimize the model using AdamW optimizer [36] with a learning rate of $1\times{1}0^{-3}$, for a total training iteration of $2\times{10}^{5}$. Validation is performed in every $2\times{10}^{3}$ iterations. We set the value of $T=100$ and ${\beta}$ value increases linearly from from $1\times{10}^{-4}$ to $6\times{10}^{-2}$ over the diffusion steps. While the DiffSinger paper [1] describes the computation of the shallow diffusion step $q$ using boundary prediction mechanism, the official GitHub implementation [30] sets $q=60$, which value we have retained in our implementation.
@@
The LAPS-Diff experiments are implemented in python 3.8.20 using pytorch 2.4.1 and numpy 1.19.2. Librosa 0.8.0 is used for audio preprocessing. All experiments run on an NVIDIA A100 GPU with 80GB memory. The training process takes approximately 2 GPU days to complete for the proposed model with the major component for complexity coming from the extra pitch calculation at each training iteration from the JDCNet [19] model.
@@ V Results and Discussion
The performance of the proposed LAPS-Diff model is compared with the SOTA DiffSinger model using both objective and subjective tests, on our Hindi Bollywood SVS dataset. Additionally, an ablation study is carried out to demonstrate the effectiveness of the individual components integrated into LAPS-Diff to adapt it for Hindi singing data, using the same evaluation methods. For objective evaluation, we have used Mel-Cepstral Distortion (MCD), logarithmic Fundamental Frequency Root Mean Square Error (logF0 RMSE), Mean Absolute Error (MAE), cosine similarity and voiced/unvoiced (V/UV) ratio [37, 38, 39] as evaluation metrics. Additionally, we have utilized the metas’s Audiobox Aesthetic Score (AES) [40] for perceptual evaluation. Whereas, for subjective evaluation, we have considered Mean Opinion Score (MOS) [41].
@@ V-A Objective Evaluation
The objective evaluation metrics considered here serve distinct purposes: MCD captures timbral (spectral) similarity, logF0 RMSE evaluates pitch similarity, MAE measures overall (global) structural similarity between mel-spectrograms, cosine similarity (with ECAPA-TDNN embeddings [42]) assesses the similarity of speaker dependent characteristics, and the V/UV ratio reflects the model’s ability to distinguish voiced and unvoiced regions. In contrast, AES provides a perceptual score using Meta’s pre-trained audio toolbox. Collectively, these metrics provide a comprehensive assessment of the generated samples across key acoustic dimensions.
@@
Meta’s AES evaluation metrics [40], adopted in our work, are designed to assess the aesthetic quality of diverse audio types including speech, music and general sounds, using a pre-trained model that closely aligns with human perceptual judgments. We select the measures relevant to synthesis from music score (i.e. content is pre-defined), viz. Production Quality (PQ) capturing the technical clarity and fidelity of the audio, including its dynamics, frequency balance, and spatial attributes. Content Enjoyment (CE) evaluates the emotional and artistic expressiveness of the audio—highlighting aspects like creativity, mood, and subjective appeal. Since we have a single voice, the AES ”production complexity” measure (PC) is irrelevant.
@@ V-A1 Visual Analysis
To compare the lyrical content preservation capabilities of the proposed model and the baseline, we extracted Hindi content embeddings using IndicWav2Vec from LAPS-Diff, DiffSinger, and the ground truth audio. These high-dimensional embeddings are projected onto a 2D space using Principal Component Analysis (PCA) [43] for visualization. As illustrated in Fig. 3, the LAPS-Diff embeddings are located closer to the ground truth than those from DiffSinger, indicating that LAPS-Diff captures content information more effectively.
@@ V-B1 MOS Testing
We conducted MOS subjective evaluation test to assess the naturalness of the generated audio samples and the ground truth. A total of 15 listeners, aged between 20-35 years and without any known hearing impairments, participated. Each participant was asked to rate the samples on a 5-point scale [41]. Our test set consists of 4 songs: one each with fast and slow singing rates and two with average singing rates. The evaluation included $6$ segments (out of a total of 28) from our test set, ensuring that we take segments from each category of fast, slow, and average singing rate. We take segments from each of the following $7$ categories: ground truth, DiffSinger, LAPS-Diff (proposed model), and four ablation variants, resulting in a total of 42 samples. These were randomly permuted and arranged in $6$ test sets of $7$ each. The results of the MOS subjective evaluation test are shown in Table III.

### v2

@@ III-D Inference Stage
In the inference stage, the optimal auxiliary decoder $AD^{*}(\cdot)$ generates the mel-spectrogram $\hat{M}$, which contains nuanced representations of the lyrical content, vocal style, and prosody. The pre-trained HiFi-GAN vocoder [Ref33:HiFi-GAN] is then used to reconstruct the audio waveform from $\hat{M}$, from which the embeddings $\boldsymbol{e}_{{mert}}$ and $\boldsymbol{e}_{{vec}}$ are extracted. These embeddings along with $\boldsymbol{e}_{{c}}$, $\boldsymbol{e}_{{t}}$, and the intermediate boundary-predicted [Ref1-Diffsinger] representation $\hat{M}_{q}$ are used to condition the optimal denoiser $D^{*}(\cdot)$, as illustrated in Fig. 2 to generate the final output. Here, $\hat{M}_{q}$ is derived from $\hat{M}$ through the shallow diffusion mechanism of DiffSinger [Ref1-Diffsinger], where $q$ denotes the shallow diffusion step. Boundary prediction mechanism is applied only at the first iteration of the reverse diffusion, rest are same as Eq.(III-C).
@@
[…]
@@
[…]
@@
40:   $\begin{aligned} M_{t-1}={}&\frac{1}{\sqrt{\alpha_{t}}}\Bigg(M_{t}-\frac{1-\alpha_{t}}{\sqrt{1-\bar{\alpha}_{t}}}\hat{\epsilon}_{\theta}(M_{t},\boldsymbol{e}_{c},\boldsymbol{e}_{t},\\ &\boldsymbol{e}_{\text{mert}},\boldsymbol{e}_{\text{vec}})\Bigg)+\sigma_{t}z\end{aligned}$
@@ IV-A Data Preprocessing
The overall duration of the singing voice segments of our Hindi Bollywood SVS dataset, introduced in Section II, is approximately 65 minutes. With 38 unique songs, segmented into 397 sung phrases each with duration ranging from 5 to 15 seconds. The dataset splits are based on distributing the songs across training, validation, and test sets with details shown in Table I. Each audio file is sampled at $16$ kHz with $16$-bit quantization. The Hindi text (Devanagari script) is converted into its corresponding phonemes using the IIT-M Hindi phoneset [Ref33:IIT-M-Phoneset]. The size of the phoneme vocabulary considered is $43$. We extract mel-spectrograms considering $80$ mel-bins using frame size of $512$ and hop size of $128$. Similar to DiffSinger [Ref1-Diffsinger], the mel-spectrograms are normalized to the range [1,1][-1,1].
@@ IV-A1 Implementation Details
In our work, we use the IndicBERT model [Ref11-IndicBERT], pre-trained on $12$ Indian languages, including Hindi. In contrast, the pre-trained XPhoneBERT model [Ref12-XPhoneBERT], which follows the BERT-base architecture [Ref33-Summation-of-embeddings], is trained on nearly $100$ languages, also including Hindi. The embedding dimension for both IndicBERT and XPhoneBERT is $768$. Linear projection is applied to reduce the dimensionality to 256, ensuring compatibility with the $256$-dimensional music score embedding $\boldsymbol{e}_{m}$. We retain the encoder architecture $E(.)$ as used in DiffSinger [DiffSinger-code].
@@
The style encoder $S(.)$ consists of four 2D residual blocks [Ref13-StyleTTS], and the style vector dimensionality is set to 48. We use the checkpoint of the pre-trained JDCNet (a ConvBi-LSTM model consisting of three ResNet blocks) [Ref15-JDCNet] from StyleTTS2 [Ref14-StyleTTS2] to extract pitch values (the pitch sequence is of variable length depending on the segment duration). Unlike the loss functions described in the original DiffSinger paper [Ref1-Diffsinger], the official GitHub implementation [DiffSinger-code] incorporates three additional duration losses—word, phone, and slur—duration along with a pitch and energy loss, employing duration, pitch and energy predictors similar to FastSpeech 2 [FastSpeech_2]. These additional losses are also used to train the auxiliary decoder of both DiffSinger and our proposed model in all the experiments. The model architecture of $AD(.)$ is the same as FastSpeech2 [Ref32-FastSpeech2, Ref1-Diffsinger]). In LAPS-Diff, the denoiser architecture $D(.)$ follows a non-causal WaveNet framework [WaveNet], similar to that used in DiffSinger [Ref1-Diffsinger].
@@
The MERT model considered in our work is trained on diverse datasets comprising musical instruments and singing voices [Ref17-MERT]. We use the pre-trained model to obtain the musical embeddings of size 1024. The IndicWav2Vec [Ref18-Indicwav2vec] is trained on nine Indian languages including Hindi, is capable of extracting linguistic content from speech audio. We employ the pre-trained IndicWav2Vec to obtain content embeddings of size 1024. We consider batch size as 48 and optimize the model using AdamW optimizer [adamw] with a learning rate of $1\times{1}0^{-3}$, for a total training iteration of $2\times{10}^{5}$. Validation is performed in every $2\times{10}^{3}$ iterations. We set the value of $T=100$ and $\beta$} value increases linearly from from $1\times{10}^{-4}$ to $6\times{10}^{-2}$ over the diffusion steps. While the DiffSinger paper [Ref1-Diffsinger] describes the computation of the shallow diffusion step $q$ using boundary prediction mechanism, the official GitHub implementation [DiffSinger-code] sets $q=60$, which value we have retained in our implementation.
@@
The LAPS-Diff experiments are implemented in python 3.8.20 using pytorch 2.4.1 and numpy 1.19.2. Librosa 0.8.0 is used for audio preprocessing. All experiments run on an NVIDIA A100 GPU with 80GB memory. The training process takes approximately 2 GPU days to complete for the proposed model with the major component for complexity coming from the extra pitch calculation at each training iteration from the JDCNet [Ref15-JDCNet] model.
@@ V Results and Discussion
The performance of the proposed LAPS-Diff model is compared with the SOTA DiffSinger model using both objective and subjective tests, on our Hindi Bollywood SVS dataset. Additionally, an ablation study is carried out to demonstrate the effectiveness of the individual components integrated into LAPS-Diff to adapt it for Hindi singing data, using the same evaluation methods. For objective evaluation, we have used Mel-Cepstral Distortion (MCD), logarithmic Fundamental Frequency Root Mean Square Error (logF0 RMSE), Mean Absolute Error (MAE), cosine similarity and voiced/unvoiced (V/UV) ratio [Ref28-MCD, ECAPATDNN-Speaker-Similarity, Ref30-VUV-Ratio] as evaluation metrics. Additionally, we have utilized the metas’s Audiobox Aesthetic Score (AES) [Ref26-AAT] for perceptual evaluation. Whereas, for subjective evaluation, we have considered Mean Opinion Score (MOS) [Ref27-MOS].
@@ V-A Objective Evaluation
The objective evaluation metrics considered here serve distinct purposes: MCD captures timbral (spectral) similarity, logF0 RMSE evaluates pitch similarity, MAE measures overall (global) structural similarity between mel-spectrograms, cosine similarity (with ECAPA-TDNN embeddings [ECAPA-TDNN]) assesses the similarity of speaker dependent characteristics, and the V/UV ratio reflects the model’s ability to distinguish voiced and unvoiced regions. In contrast, AES provides a perceptual score using Meta’s pre-trained audio toolbox. Collectively, these metrics provide a comprehensive assessment of the generated samples across key acoustic dimensions.
@@
Meta’s AES evaluation metrics [Ref26-AAT], adopted in our work, are designed to assess the aesthetic quality of diverse audio types including speech, music and general sounds, using a pre-trained model that closely aligns with human perceptual judgments. We select the measures relevant to synthesis from music score (i.e. content is pre-defined), viz. Production Quality (PQ) capturing the technical clarity and fidelity of the audio, including its dynamics, frequency balance, and spatial attributes. Content Enjoyment (CE) evaluates the emotional and artistic expressiveness of the audio—highlighting aspects like creativity, mood, and subjective appeal. Since we have a single voice, the AES ”production complexity” measure (PC) is irrelevant.
@@ V-A1 Visual Analysis
To compare the lyrical content preservation capabilities of the proposed model and the baseline, we extracted Hindi content embeddings using IndicWav2Vec from LAPS-Diff, DiffSinger, and the ground truth audio. These high-dimensional embeddings are projected onto a 2D space using Principal Component Analysis (PCA) [PCA] for visualization. As illustrated in Fig. 3, the LAPS-Diff embeddings are located closer to the ground truth than those from DiffSinger, indicating that LAPS-Diff captures content information more effectively.
@@ V-B1 MOS Testing
We conducted MOS subjective evaluation test to assess the naturalness of the generated audio samples and the ground truth. A total of 15 listeners, aged between 20-35 years and without any known hearing impairments, participated. Each participant was asked to rate the samples on a 5-point scale [Ref27-MOS]. Our test set consists of 4 songs: one each with fast and slow singing rates and two with average singing rates. The evaluation included $6$ segments (out of a total of 28) from our test set, ensuring that we take segments from each category of fast, slow, and average singing rate. We take segments from each of the following $7$ categories: ground truth, DiffSinger, LAPS-Diff (proposed model), and four ablation variants, resulting in a total of 42 samples. These were randomly permuted and arranged in $6$ test sets of $7$ each. The results of the MOS subjective evaluation test are shown in Table III.